# Minimal Fractional Topological Insulator in half-filled conjugate moiré Chern bands

### v2

@@ IV Bootstrapping the minimal FTI
In this section, we bootstrap the minimal topological order (TO) allowed by the Hall responses under the charge and the $S_{z}$-spin U$(1)$ symmetries and show that ${\cal D}^{2}=32$ is the minimal squared total quantum dimension compatible with a spin-Hall conductivity $\sigma^{\rm sh}=1/2$, zero electric Hall responses, and zero Hall response with respect to the spin $S_{z}$ quantum number. The vanishing of the diagonal Hall responses associated with either the electric charge or $S_{z}$ is required by the time-reversal symmetry.
For technical convenience, we prefer to re-organize the charge and spin U(1) symmetries into $\mathrm{U}(1)_{\uparrow}\times$ U(1), where $\mathrm{U}(1)_{\uparrow,\downarrow}$ are associated with the charge conversation within each of the valleys 1 and 2 (which are locked to electron spins $\uparrow$ and $\downarrow$). The Hall responses we focus on in this paper are equivalent to the combination of (1) $\mathrm{U}(1)_{\uparrow}$-Hall conductivity $\sigma^{\rm h}_{1}=1/2$, (2) $\mathrm{U}(1)_{\downarrow}$-Hall conductivity $\sigma^{\rm h}_{2}=-1/2$, and (3) zero mixed $\mathrm{U}(1)_{\uparrow}\times\mathrm{U}(1)_{\downarrow}$ response. For an anyon $x$, we denote its fractional charge under $\mathrm{U}(1)_{\uparrow}\times\mathrm{U}(1)_{\downarrow}$ as $q_{x}=(q_{x,1},q_{x,2})$, which is defined modulo the charge of local bosonic objects, i.e. $q_{x}\sim q_{x}+(n,m)$ for $n,m\in\mathbb{Z}$ with $m+n$ even. The fractional electric charge and $S_{z}$-spin of the anyon $x$ is then given by $(q_{x,1}+q_{x,2})e$ and $(q_{x,1}-q_{x,2})\hbar/2$. In the TO, the electrons and holes are all identified as the transparent fermionic particle $f$ with $q_{f}=(\pm 1,0)\sim(0,\pm 1)$. Two such $f$ particles fuse into a trivial anyon $\mathbf{1}$, i.e. $f\times f={\bf 1}$
@@
For minimality, ${\cal C}_{\rm I}$ should be the entire set of anyons, and each anyon should be Abelian, which yields a squared total quantum dimension ${\cal D}^{2}=32$. The $\mathbb{Z}_{4}$ FTI constructed in Sec. II provides an example of such type-I minimal TOs. $v_{1}$, $v_{2}$, and $y$ correspond to the anyons $\mathbf{e}\mathbf{m}$ and $\mathbf{e}^{3}\mathbf{m}$, $\mathbf{e}$ respectively. Note that the $\mathbb{Z}_{4}$ FTI not only has the desired Hall responses but also preserves the time-reversal symmetry. Moreover, ${\bf m}^{2}$, which corresponds to $v_{1}v_{2}$, is the anyon generated by the insertion of a $2\pi$ flux in each valley. Its time-reversal conjugate should be associated with a $-2\pi$ flux in each valley, which still belongs to the same anyon type ${\bf m}^{2}$ given that $v_{1}^{2}v_{2}^{2}\sim{\bf 1}$. Knowing that ${\bf m}^{2}$ is related to its time-reversal partner by a $4\pi$ flux in each valley and that the spin Hall conductivity is $\sigma^{\rm sh}=1/2$, ${\bf m}^{2}$ must be a Kramers doublet under time reversal Levin and Stern (2009), which is consistent with the analysis in Sec. III. This statement has used the fact that the system is ultimately built out of the electron and holes that are Kramers doublets with ${\cal T}^{2}=-1$.
@@
We emphasize that the bootstrap of type-I and type-II minimal TOs only uses charge conservation, spin-$S_{z}$ conservation, and the Hall transport signals. In the Appendix, we further show that once time-reversal symmetry ${\cal T}$ is considered, the $\mathbb{Z}_{4}$ FTI constructed in Sec. II is the unique mFTI with the minimal total quantum dimension.
@@ V.1 Descending from larger topological orders
The anyon $(4,4)$ is obviously a self-boson, hence it can condense. The anyon $(4,4)$ can be created by adiabatically inserting $4\pi$ fluxes of $\mathrm{U}(1)_{\uparrow}$ and $\mathrm{U}(1)_{\downarrow}$ in both valleys. It carries spin $S_{z}=\hbar$. To avoid spontaneously breaking the spin $S_{z}$ symmetry in the condensate, we bind the $(4,4)$ anyon with the physical inter-valley exciton $b^{s}$ with $S_{z}=-\hbar$ to neutralize its $S_{z}$ spin. The mutual braiding statistical between the neutralized bosonic anyon $(4,4)$ and a general anyon $(n_{1},n_{2})$ is $e^{\mathrm{i}\pi(n_{1}-n_{2})}$. Hence, the condensate of $(4,4)$ would confine all the anyons whose $n_{1}$ and $n_{2}$ have opposite parity. The condensate also identifies anyons $(n_{1},n_{2})$ and $(n_{1}+4,n_{2}+4)$. Consequently, there are, in total, 16 anyons left (before including the electrons/holes), the same as the mFTI. Using the notation of Sec. IV, the anyon $v_{1}$ and $v_{2}$ correspond to $(2,0)$, and $(0,2)$. One of the most elementary deconfined anyons is $(1,1)$, which is a self-boson with electric charge 00 and spin $S_{z}=\hbar/4$. Hence, $(1,1)$ is identified with the anyon $\bf{m}$ of the mFTI. Similarly, the anyon $(1,-1)$ can be identified as anyon $\bf{e}$ of the mFTI. The anyons $(1,1)$ and $(1,-1)$ also have the desired mutual braiding statistics $e^{\mathrm{i}\pi/2}$. At this point, we can confirm that the condensate of the neutralized anyon $(4,4)$ is the mFTI with $\mathbb{Z}_{4}$ TO. As an alternative approach, how the $\mathrm{U}(1)_{8}\times\mathrm{U}(1)_{-8}$ TO descends, via the anyon condensation of (4,4), to the desired mFTI can also be shown explicitly using the $K$-matrix formalism.
@@
In our current case, the mFTI can be constructed through the same procedure. The SPT state we start with could be constructed by the inter-valley exciton operator $b^{s}$, and “one quarter” of the inter-valley Cooper pair $b^{c}$. In other words, we formally fractionalize $b^{c}$ into four bosonic partons each carrying charge-$e/2$: $b^{c}=(\tilde{b}^{c})^{4}$. Then, $\tilde{b}^{c}$ must coupled to a $Z_{4}$ gauge field. Now, we construct a minimal level-1 bosonic SPT state between $b^{s}$ and $\tilde{b}_{c}$ Senthil and Levin (2013), which can be described in various formulations, including the Chern-Simons theory Lu and Vishwanath (2012), or the nonlinear sigma model Senthil and Levin (2013); Xu and Senthil (2013). The most convenient formulation for our current purpose is the Chern-Simons theory developed in Ref. Lu and Vishwanath, 2012. The Chern-Simons theory for many $(2+1)d$ bosonic SPT states takes a universal form:
@@ V.3 Example of type-II minimal TO
This non-Abelian TO, described by the condensate of $(\mathds{1},4,4)$, is one of the type-II minimal TO that can be realized in a pair of half-filled conjugate Chern bands. It has the same Hall responses with respect to the electric charge and the $S_{z}$ quantum number as the mFTI. However, this state breaks the time-reversal symmetry because it has the same chiral central charge $c=1/2$ as the original $\mathrm{Pfaffian}\times\mathrm{U}(1)_{-8}$ TO. Consequently, this type-II minimal TO also has a quantized thermal-Hall response.
@@
Now, we generalize the study of mFTI to a pair of conjugate Chern bands with $1/q$ electron filling in each band. Here, $q$ is a positive integer. It is natural to expect such a mFTI to have a fractional spin-Hall conductivity $\sigma^{\rm sh}=1/q$ and vanishing diagonal Hall responses for the electric charge and the spin-$S_{z}$ quantum number. The symmetry of such a mFTI includes the charge conservation, $S_{z}$ conservation, and time-reversal symmetry ${\cal T}$. Per our definition, the mFTI is minimal in the sense that it has the smallest total quantum dimension among all the TOs with the required symmetries and Hall transport signals with the given $q$. We will separately discuss the mFTIs with even and odd $q$’s. And we will provide a classification of them.
For even $q$, by generalizing Sec. IV, we can show that there is a list of, in total, eight different minimal TOs compatible with charge conversation, $S_{z}$ conservation, and the Hall transports described above. The details are given in the Appendix. After further imposing time-reversal symmetry ${\cal T}$, the list narrows down to a unique mFTI, which has the $\mathbb{Z}_{2q}$ TO, the same TO as the $\mathbb{Z}_{2q}$ toric code. The squared total quantum dimension (with the electron included) is ${\cal D}^{2}=8q^{2}$. The minimally-charged anyon has an electric charge $e^{\ast}=e/q$, and the minimal-$S_{z}$ anyon has the $S_{z}$ quantum number $\hbar/{2q}$. Such an mFTI is not a valley-decoupled product TO. Since $\sigma^{\rm sh}/e^{\ast}$ is odd, based on the result of Ref. Levin and Stern, 2009, the gapless edge state of the mFTI is protected by charge conservation and time-reversal symmetry (even when $S_{z}$ is not conserved). The edge theory is a straightforward generalization of Eq. (3)
${\cal L}_{\mathrm{edge}}=\frac{2q}{4\pi}\mathrm{i}\partial_{\tau}\phi^{c}\partial_{x}\phi^{s}-\frac{2q}{4\pi}\mathrm{i}\partial_{\tau}\phi^{s}\partial_{x}\phi^{c}+\cdots,$ (22)
@@
$\sigma^{\rm sh}_{\rm tot}=\frac{1}{q_{2}}+\frac{1}{q_{2}}=\frac{2^{k_{1}-k_{2}}p_{1}+p_{2}}{2^{k}_{1}p_{1}p_{2}}.$ (24)
@@
VI Summary
@@
Motivated by a recent experimentKang et al. (2024), in this work, we proposed a minimal fractional topological insulator that can potentially be realized in a pair of conjugate Chern bands, both with electron filling $\nu=1/2$. The mFTI we proposed has the minimal total quantum dimension amongst all the topological orders that can be potentially realized in the same setting (especially with the same symmetry and the same experimentally observed fractionally quantized transport signal). We caution that a more thorough analysis, similar to that for the $\nu=2/3$ fractional quantum stateKane et al. (1994); Kane and Fisher (1995a, b), is required to understand the transport quantization. We found that the minimally-charged anyon in mFTI has a fractional charge $e^{\ast}=e/2$, which is different from all product TOs. The gapless edge state of this mFTI is protected by charge conservation and time-reversal symmetry even when $S_{z}$ is not conserved. We also showed that the mFTI is the common descendant of various product TOs of the two valleys, which have larger quantum dimensions. The mFTI is also the common symmetry-enriched TO that can be promoted from different bosonic SPT states by gauging the discrete subgroup of their symmetries.
@@
If the constraint on time-reversal symmetry is released (while maintaining the Hall transports requirements associated with the electric charge and $S_{z}$), we find mFTI belonging to an 8-fold classified list of minimal TOs. Four of them, albeit time-reversal broken, are non-Abelian TOs that exhibit quantized thermal Hall effects in addition to the required Hall responses associated with the charge and $S_{z}$ quantum numbers.
@@
We generalize the construction of mFTI (and the minimal TOs) to a pair of conjugate $1/q$-filled Chern bands (for positive integer $q$). We classify all the mFTIs via the stability of the gapless interfaces between them.
The authors thank Leon Balents and Matthew Fisher for very helpful discussions. C.X. is supported by the Simons foundation through the Simons investigator program. M.C. acknowledges support from NSF under award number DMR-1846109. C.-M.J. is supported by a faculty startup grant at Cornell University. While finishing this paper, we became aware of another independent work May-Mann et al. (2024) studying the edge state of FTIs that are a product of conjugate TOs from two valleys, which, as we discussed, should have a larger quantum dimension than the mFTI constructed in this paper.

### v3

@@ IV Bootstrapping the minimal FTI
In this section, we bootstrap the minimal topological order (TO) allowed by the Hall responses under the charge and the $S_{z}$-spin U$(1)$ symmetries and show that ${\cal D}^{2}=32$ is the minimal squared total quantum dimension compatible with a spin-Hall conductance $\sigma^{\rm sh}=1/2$, zero electric Hall responses, and zero Hall response with respect to the spin $S_{z}$ quantum number. The vanishing of the diagonal Hall responses associated with either the electric charge or $S_{z}$ is required by the time-reversal symmetry. As we will see, the method used in this section is fundamentally based on the internal consistency of topological orders enriched by symmetries, and it yields the structure of the minimal TOs without relying on any microscopic details of the system. Hence, this method is a bootstrap-type approach. After we bootstrap the minimal TOs, we will discuss the physical significance of the minimal total quantum dimension.
For technical convenience, we prefer to re-organize the charge and spin U(1) symmetries into $\mathrm{U}(1)_{\uparrow}\times$ U(1), where $\mathrm{U}(1)_{\uparrow,\downarrow}$ are associated with the charge conservation within each of the valleys 1 and 2 (which are locked to electron spins $\uparrow$ and $\downarrow$). The Hall responses we focus on in this paper are equivalent to the combination of (1) $\mathrm{U}(1)_{\uparrow}$-Hall conductance $\sigma^{\rm h}_{1}=1/2$, (2) $\mathrm{U}(1)_{\downarrow}$-Hall conductance $\sigma^{\rm h}_{2}=-1/2$, and (3) zero mixed $\mathrm{U}(1)_{\uparrow}\times\mathrm{U}(1)_{\downarrow}$ response. For an anyon $x$, we denote its fractional charge under $\mathrm{U}(1)_{\uparrow}\times\mathrm{U}(1)_{\downarrow}$ as $q_{x}=(q_{x,1},q_{x,2})$, which is defined modulo the charge of local bosonic objects, i.e. $q_{x}\sim q_{x}+(n,m)$ for $n,m\in\mathbb{Z}$ with $m+n$ even. The fractional electric charge and $S_{z}$-spin of the anyon $x$ is then given by $(q_{x,1}+q_{x,2})e$ and $(q_{x,1}-q_{x,2})\hbar/2$. In the TO, the electrons and holes are all identified as the transparent fermionic particle $f$ with $q_{f}=(\pm 1,0)\sim(0,\pm 1)$. Two such $f$ particles fuse into a trivial anyon $\mathbf{1}$, i.e. $f\times f={\bf 1}$
@@
For minimality, ${\cal C}_{\rm I}$ should be the entire set of anyons, and each anyon should be Abelian, which yields a squared total quantum dimension ${\cal D}^{2}=32$. The $\mathbb{Z}_{4}$ FTI constructed in Sec. II provides an example of such type-I minimal TOs. $v_{1}$, $v_{2}$, and $y$ correspond to the anyons $\mathbf{e}\mathbf{m}$ and $\mathbf{e}^{3}\mathbf{m}$, $\mathbf{e}$ respectively. Note that the $\mathbb{Z}_{4}$ FTI not only has the desired Hall responses but also preserves the time-reversal symmetry. Moreover, ${\bf m}^{2}$, which corresponds to $v_{1}v_{2}$, is the anyon generated by the insertion of a $2\pi$ flux in each valley. Its time-reversal conjugate should be associated with a $-2\pi$ flux in each valley, which still belongs to the same anyon type ${\bf m}^{2}$ given that $v_{1}^{2}v_{2}^{2}\sim{\bf 1}$. Knowing that ${\bf m}^{2}$ is related to its time-reversal partner by a $4\pi$ flux in each valley and that the spin Hall conductance is $\sigma^{\rm sh}=1/2$, ${\bf m}^{2}$ must be a Kramers doublet under time reversal Levin and Stern (2009), which is consistent with the analysis in Sec. III. This statement has used the fact that the system is ultimately built out of the electron and holes that are Kramers doublets with ${\cal T}^{2}=-1$.
@@
We emphasize that the bootstrap of type-I and type-II minimal TOs only uses charge conservation, spin-$S_{z}$ conservation, and the Hall transport signals. In the Appendix, we further show that once time-reversal symmetry ${\cal T}$ is considered, the $\mathbb{Z}_{4}$ FTI constructed in Sec. II is the unique mFTI with the minimal total quantum dimension. Moreover, the analysis above fixed the anyon statistics of type-I and type-II minimal TOs up to the topological spin $\theta_{y}$. The full classifications of all the minimal TOs and the allowed values of $\theta_{y}$ are provided in the Appendix.
Now, we comment on the physical implication of the minimal total quantum dimension. The squared total quantum dimension ${\cal D}^{2}$ of a TO is a metric of the complexity of the TO, given by the sum of the squared quantum dimensions of all the anyons in the TO. It is directly related to the topological entanglement entropy that characterizes the long-range entanglement in the two-dimensional TO.Kitaev and Preskill (2006); Levin and Wen (2006) Therefore, the search for the minimal TOs is equivalent to identifying the TOs with the least amount of long-range entanglement allowed by the symmetry and transport requirements. One may expect the minimal TOs to be, in general, more robust than the more complex ones that require more long-range entanglement. Additionally, for Abelian TOs, ${\cal D}^{2}$ is equal to the number of anyons and, consequently, the ground state degeneracy on the torus (when the electron is treated as a trivial anyon). Therefore, generically speaking, if a minimal TO is Abelian, it must also be the TO with the smallest ground state degeneracy amongst all the Abelian TOs satisfying the symmetry and transport requirements. However, we caution that the total quantum dimensions of non-Abelian TOs are different from their ground state degeneracies on the torus. In the current context of minimal TOs with spin-Hall conductance $\sigma^{\rm sh}=1/2$, we notice that the Abelian mFTI and other type-I minimal TOs have the same ground state degeneracy 16 on the torus, while the type-II minimal TOs’ ground state degeneracies are 12. Despite the difference in ground state degeneracies, all the minimal TOs share the same total quantum dimensions.
As a side note, the previous results in fractional quantum states under a magnetic field provide further justification for the general idea of bootstrapping TOs with minimal quantum dimensions that are compatible with observed transport. The celebrated $1/m$ Laughlin states (with $m$ odd) are the TOs with the minimal quantum dimensions compatible with electrical Hall conductance $1/m$. For a half-filled Landau level, all of the leading incompressible fractional quantum Hall candidates, including the Pfaffian state Moore and Read (1991); Greiter et al. (1991); Wen (1993), anti-Pfaffian state Levin et al. (2007), PH-Pfaffian state Son (2015), the Abelian $\mathrm{U}(1)_{8}$ state Greiter et al. (1992), and the “$331$” state Wen and Zee (1992); Overbosch and Wen (2008), are the minimal TOs compatible with electrically Hall conductance $1/2$ (modulo integers). Therefore, bootstrapping the minimal TOs using transport and symmetry can be a powerful tool for the study of topological quantum matter.
@@ V.1 Descending from larger topological orders
The anyon $(4,4)$ is obviously a self-boson, hence it can condense. The anyon $(4,4)$ can be created by adiabatically inserting $4\pi$ fluxes of $\mathrm{U}(1)_{\uparrow}$ and $\mathrm{U}(1)_{\downarrow}$ in both valleys. It carries spin $S_{z}=\hbar$. To avoid spontaneously breaking the spin $S_{z}$ symmetry in the condensate, we bind the $(4,4)$ anyon with the physical inter-valley exciton $b^{s}$ with $S_{z}=-\hbar$ to neutralize its $S_{z}$ spin. The mutual braiding statistics between the neutralized bosonic anyon $(4,4)$ and a general anyon $(n_{1},n_{2})$ is $e^{\mathrm{i}\pi(n_{1}-n_{2})}$. Hence, the condensate of $(4,4)$ would confine all the anyons whose $n_{1}$ and $n_{2}$ have opposite parity. The condensate also identifies anyons $(n_{1},n_{2})$ and $(n_{1}+4,n_{2}+4)$. Consequently, there are, in total, 16 anyons left (before including the electrons/holes), the same as the mFTI. Using the notation of Sec. IV, the anyon $v_{1}$ and $v_{2}$ correspond to $(2,0)$, and $(0,2)$. One of the most elementary deconfined anyons is $(1,1)$, which is a self-boson with electric charge 00 and spin $S_{z}=\hbar/4$. Hence, $(1,1)$ is identified with the anyon $\bf{m}$ of the mFTI. Similarly, the anyon $(1,-1)$ can be identified as anyon $\bf{e}$ of the mFTI. The anyons $(1,1)$ and $(1,-1)$ also have the desired mutual braiding statistics $e^{\mathrm{i}\pi/2}$. At this point, we can confirm that the condensate of the neutralized anyon $(4,4)$ is the mFTI with $\mathbb{Z}_{4}$ TO. As an alternative approach, how the $\mathrm{U}(1)_{8}\times\mathrm{U}(1)_{-8}$ TO descends, via the anyon condensation of (4,4), to the desired mFTI can also be shown explicitly using the $K$-matrix formalism.
@@
In our current case, the mFTI can be constructed through the same procedure, as indicated in Fig. 3. The SPT state we start with could be constructed by the inter-valley exciton operator $b^{s}$, and “one quarter” of the inter-valley Cooper pair $b^{c}$. In other words, we formally fractionalize $b^{c}$ into four bosonic partons each carrying charge-$e/2$: $b^{c}=(\tilde{b}^{c})^{4}$. Then, $\tilde{b}^{c}$ must coupled to a $Z_{4}$ gauge field. Now, we construct a minimal level-1 bosonic SPT state between $b^{s}$ and $\tilde{b}_{c}$ Senthil and Levin (2013), which can be described in various formulations, including the Chern-Simons theory Lu and Vishwanath (2012), or the nonlinear sigma model Senthil and Levin (2013); Xu and Senthil (2013). The most convenient formulation for our current purpose is the Chern-Simons theory developed in Ref. Lu and Vishwanath, 2012. The Chern-Simons theory for many $(2+1)d$ bosonic SPT states takes a universal form:
@@ V.3 Example of type-II minimal TO
This non-Abelian TO, described by the condensate of $(\mathds{1},4,4)$, is one of the type-II minimal TO that can be realized in a pair of half-filled conjugate Chern bands. It has the same Hall responses with respect to the electric charge and the $S_{z}$ quantum number as the mFTI. However, this state breaks the time-reversal symmetry because it has the same chiral central charge $c=1/2$ as the original $\mathrm{Pfaffian}\times\mathrm{U}(1)_{-8}$ TO. Consequently, this type-II minimal TO also has a quantized thermal-Hall response given by its chiral central charge $c=1/2$.
@@
Now, we generalize the study of mFTI to a pair of conjugate Chern bands with $1/q$ electron filling in each band. Here, $q$ is a positive integer. It is natural to expect such a mFTI to have a fractional spin-Hall conductance $\sigma^{\rm sh}=1/q$ and vanishing diagonal Hall responses for the electric charge and the spin-$S_{z}$ quantum number. The symmetry of such a mFTI includes the charge conservation, $S_{z}$ conservation, and time-reversal symmetry ${\cal T}$. Per our definition, the mFTI is minimal in the sense that it has the smallest total quantum dimension among all the TOs with the required symmetries and Hall transport signals with the given $q$. We will separately discuss the mFTIs with even and odd $q$’s. And we will provide a classification of them.
For even $q$, by generalizing Sec. IV, we can show that there is a list of, in total, eight different minimal TOs compatible with charge conservation, $S_{z}$ conservation, and the Hall transports described above. The details are given in the Appendix. After further imposing time-reversal symmetry ${\cal T}$, the list narrows down to a unique mFTI, which has the $\mathbb{Z}_{2q}$ TO, the same TO as the $\mathbb{Z}_{2q}$ toric code. The squared total quantum dimension (with the electron included) is ${\cal D}^{2}=8q^{2}$. The minimally-charged anyon has an electric charge $e^{\ast}=e/q$, and the minimal-$S_{z}$ anyon has the $S_{z}$ quantum number $\hbar/{2q}$. Such an mFTI is not a valley-decoupled product TO. Since $\sigma^{\rm sh}/e^{\ast}$ is odd, based on the result of Ref. Levin and Stern, 2009, the gapless edge state of the mFTI is protected by charge conservation and time-reversal symmetry (even when $S_{z}$ is not conserved). The edge theory is a straightforward generalization of Eq. (3)
${\cal L}_{\mathrm{edge}}=\frac{2q}{4\pi}\mathrm{i}\partial_{\tau}\phi^{c}\partial_{x}\phi^{s}+\frac{2q}{4\pi}\mathrm{i}\partial_{\tau}\phi^{s}\partial_{x}\phi^{c}+\cdots,$ (23)
@@
$\sigma^{\rm sh}_{\rm tot}=\frac{1}{q_{1}}+\frac{1}{q_{2}}=\frac{2^{k_{1}-k_{2}}p_{1}+p_{2}}{2^{k_{1}}p_{1}p_{2}}.$ (25)
@@
VI Summary and Discussion
Motivated by a recent discovery of the fractional quantum spin Hall effect in twisted bilayer MoTe2Kang et al. (2024), in this work, we proposed a minimal fractional topological insulator that can potentially be realized in a pair of conjugate Chern bands, both at electron filling $\nu=1/2$. The mFTI we proposed has the minimal total quantum dimension amongst all the topological orders that can be potentially realized in the same setting (especially with the same symmetry and the fractionally quantized spin-Hall conductance $\sigma^{\rm sh}=1/2$ suggested by the experiment on twisted bilayer MoTe2).
@@ VI Summary and Discussion
We found that the minimally-charged anyon in mFTI has a fractional charge $e^{\ast}=e/2$, which is different from all product TOs. The gapless edge state of this mFTI is protected by charge conservation and time-reversal symmetry even when $S_{z}$ is not conserved. We also showed that the mFTI is the common descendant of various product TOs of the two valleys, which have larger quantum dimensions. The mFTI is also the common symmetry-enriched TO that can be promoted from different bosonic SPT states by gauging the discrete subgroup of their symmetries.
We caution that a more thorough analysis, similar to that for the $\nu=2/3$ fractional quantum stateKane et al. (1994); Kane and Fisher (1995a, b), is required to understand the quantization of the transport signals measured in the 4-terminal geometry considered in experiment Ref. Kang et al., 2024. In particular, to compare our mFTI state with this experiment, we need to carefully examine the interactions between the edge modes, disorder, and the coupling between the edge modes and the electrical contact, which would be an interesting topic for future studies.
@@
When the constraint on time-reversal symmetry is released (while maintaining the Hall transports requirements associated with the electric charge and $S_{z}$), we found mFTI belonging to an 8-fold classified list of minimal TOs. The mFTI and 3 other minimal TOs are Abelian topological orders, while the remaining 4 are non-Abelian. All 7 minimal TOs other than the mFTI exhibit quantized thermal Hall effects in addition to the required Hall responses associated with the charge and $S_{z}$ quantum numbers. In particular, it is worth noting that the simplest non-Abelian minimal TOs must break the time-reversal symmetries.
@@
We generalized the construction of mFTI (and the minimal TOs) to a pair of conjugate $1/q$-filled Chern bands (for positive integer $q$). We classify all the mFTIs via the stability of the gapless interfaces between them.
It is interesting to explore the mFTI’s and the minimal TOs’ nearby phases and transitions in future research. For example, condensing the ${\bf e}$ anyon in the mFTI can result in superconductivity, while the condensation of ${\bf m}$ can result in a state with a spontaneously broken $S_{z}$ symmetry, equivalent to an inter-valley coherent state (due to spin-valley locking in the moiré TMD systems). It would be interesting to investigate the nature of these transitions and the details of the resulting phases. We note that recently both a superconductor Xia et al. (2024); Guo et al. (2024) and an insulator Xia et al. (2024) were observed in a pair of conjugate Chern bands at half-filling, similar to our set-up, and the insulator phase may indeed have inter-valley coherent spin order Devakul et al. (2021b). It would be interesting to explore the detailed relation between the mFTI discussed in this work and the observed superconductivity and insulating phase. Another direction is to explore possible compressible phases in the vicinity of the mFTI and minimal TOs. Examples of such compressible phases include the conjugate composite Fermi liquid Myerson-Jain et al. (2023) and the vortex spin liquid Zhang (2024a). Investigating the universal properties of the transitions between the compressible states and the minimal TOs would be an interesting topic for future studies.
In this work, we’ve identified the minimal TOs under the symmetry and spin-Hall transport requirements. They have the smallest possible total quantum dimensions. And they are different from the product TOs. It would also be interesting to investigate if there are other possible topological state candidates satisfying the same symmetry and transport requirements that are simpler, namely smaller in the total quantum dimensions, than the product TOs.
Acknowledgement - The authors thank Leon Balents and Matthew Fisher for very helpful discussions. C.X. is supported by the Simons foundation through the Simons investigator program. M. C. acknowledges support from NSF under Award No. DMR-1846109, and is currently supported in part by NSF under Award No. DMR-2424315. C.-M.J. is supported by the Alfred P. Sloan Foundation through a Sloan Research Fellowship.
While finishing this paper, we became aware of another independent work May-Mann et al. (2024) studying the edge state of FTIs that are a product of conjugate TOs from two valleys, which, as we discussed, should have a larger quantum dimension than the mFTI constructed in this paper. Shortly after our work was posted on Arxiv, an independent workZhang (2024b) studying the possible $p$-wave pairings on the neutral Fermi surface of the vortex spin liquid appeared. The resulting phases turn out to be topologically equivalent to our $\mathbb{Z}_{4}$ mFTI and the non-Abelian minimal TO discussed in Sec. V.3.